OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1,-1) and (5, 3)? (Write it in slope intercept form: y = mx +b)

OpenStudy (anonymous):

y = mx+b Where: m is the slope, and b is the y-intercept...

OpenStudy (anonymous):

let's find what m is, the slope of the line...

OpenStudy (anonymous):

For lines like these, the slope is always defined as "the change in y over the change in x" or, in equation form: So we use the slope intercept formula to solve...

OpenStudy (anonymous):

(y2−y1)(x2−x1)

OpenStudy (anonymous):

plug in the numbers into the formula. m=3- (-1)\5- (-1) which would be 3+1/5+1= 4/6

OpenStudy (anonymous):

=2/3 simplified.

OpenStudy (anonymous):

So, we have the first piece to finding the equation of this line, and we can fill it into y=mx+b like this:

OpenStudy (anonymous):

y=2/3x+b

OpenStudy (anonymous):

To find b, think about what your (x,y) points mean: (-1,-1). When x of the line is -1, y of the line must be -1. (5,3). When x of the line is 5, y of the line must be 3.

OpenStudy (anonymous):

You can use either (x,y) point you want..the answer will be the same: (-1,-1). y=mx+b or -1=2/3 × -1+b, or solving for b: b=-1-(2/3)(-1). b=-1/3.

OpenStudy (anonymous):

(5,3). y=mx+b or 3=2/3 × 5+b, or solving for b: b=3-(2/3)(5). b=-1/3.

OpenStudy (anonymous):

The equation of the line that passes through the points (-1,-1) and (5,3) is y=2/3x-1/3